# PoIFusion: Multi-Modal 3D Object Detection via Fusion at Points of Interest

### v1

@@ 4.2 Comparison with Other Methods
Table 2: Comparison of 3D multi-object tracking performance. The results are evaluated on the nuScenes validation set.
Method mAP NDS Latency (ms)
A100
V100
Transfusion [2]
67.5
71.3
153.8
190.3
BEVFusion (MIT) [42]
68.5
71.4
113.9 135.5
DeepInteraction [75]
69.9
72.6
204.1
344.8
CMT [69]
70.3
72.9
210.8
387.1
${}^{\dagger}$ [69]
70.3
72.9
159.7
PoIFusion (ours)
71.2 73.2
115.2
163.7
@@
Table 3: Runtime comparison.normal-$\dagger$” means this model is accelerated with the Flash-Attention Operator [15]. NVIDIA V100 GPU is not compatible with the Flash Attention package.
@@
NDS
mAP
center only
72.8
70.6
center + corner
73.2 71.2
NDS
mAP
baseline
72.4
70.1
+ B.T.
72.8
70.8
+ B.T. & P.S.
73.2 71.2
NDS
mAP
static fusion
71.8
69.3
dynamic fusion
73.2 71.2
NDS
mAP
camera only
40.6
29.4
LiDAR only
70.2
65.3
camera + LiDAR
73.2 71.2
image resolution
NDS
mAP
800$\times$448
73.2
71.2
1600$\times$640
73.7 71.9
voxel size
NDS
mAP
(0.125, 0.125, 0.2)
71.6
69.1
(0.1, 0.1, 0.2)
72.3
69.8
(0.075, 0.075, 0.2)
73.2 71.2
Table 4: Ablative experiments. In these experiments, the backbone networks are ResNet-50 and VoxelNet. Unless specified, image resolution is set as $800\times 448$, and the voxel size of point clouds is ($0.075m$,$0.075m$,$0.2m$). Our default setting is highlighted in gray.
Our proposed PoIFusion has set a state-of-the-art record on the nuScenes dataset. Specifically, equipped with ResNet-50, our method achieves 73.2% NDS / 71.2% mAP on the validation set and achieves 74.3% NDS / 72.6% mAP on the test set. Upgrading the image backbone to a more powerful Swin-T [41], the performance on the test set boosts to 74.9% NDS and 73.4% mAP, which makes an absolute improvement of 0.8% NDS and 1.4% mAP over the previous best results [69]. Compared to the unified-view methods [34, 42, 29], our PoIFusion better keeps modal-specific information and thus outperforms the representative BEVFusion [42] by 2.0% NDS and 3.2% mAP on the nuScenes test set. In comparison with recent work, i.e., ObjectFusion [5], which also generates object-centric features, our method performs multi-modal fusion at adaptive PoIs, instead of fusing the regional features extracted from different representation spaces. The employment of PoIs improves the flexibility of sampling location and enables fine-grained feature fusion, thus working better. Moreover, the qualitative results are shown in Figure 4, with one example from a sunny day and another example from a rainy night. Our PoIFusion can robustly perform 3D object detection under varying weather and lighting conditions.
Extended Results of MOT.
In addition to the main results of 3D object detection, we evaluate the generalizability of our method on the 3D multiple object tracking task of the nuScenes dataset. Specifically, we follow CenterPoint [76] to adopt the “tracking-by-detection” scheme that offline links the detection boxes into tracking tubelets. Note that other compared methods, except VoxelNeXt [11], also follow this same scheme. As shown in Table 2, our PoIFusion works the best (75.1% AMOTA and 50.7%AMOTP) among the compared methods, demonstrating our better detection results can also benefit other downstream tasks.
Runtime Comparison.
Furthermore, we conduct runtime comparisons among the recent works. For a fair comparison, we re-evaluate the latency on an NVIDIA A100 and an NVIDIA V100 GPU. Since the implementation of the voxelization operation is a confounder for runtime evaluation but will not influence the detection result, we exclude its time cost in comparison. As shown in Table 3, the latency of our PoIFusion is comparable to the fastest method, i.e., BEVFusion (MIT), which further shows the potential of our method to serve as a strong baseline for future investigation.
4.3 Ablative Experiments
Effect of Components in the Multi-Modal Decoder.
In this section, we analyze the design of the proposed multi-modal decoder from three aspects: (a) anchor points, (b) operations to generate PoIs, and (c) the fusion block. As shown in Table LABEL:tab:ablation:anchor, the model that only uses the center point achieves 72.8% NDS and 70.6% mAP. By additionally involving the corner points, the NDS and mAP boosts to 73.2% and 71.2%, respectively. This comparison shows the benefits of representing the geometric property of query boxes. In Table LABEL:tab:ablation:transformation, we validate the operations to derive PoIs from anchor points. The baseline directly takes the anchor points as PoIs without adaptation, achieving 70.1% mAP. By involving the holistic box-level transformation and further applying the point-wise shift, the performance boosts to 70.8 % mAP and 71.2% mAP, respectively. The online adjustment of PoIs ameliorates the misalignment in feature sampling and thereby boosts the detection result.
Table LABEL:tab:ablation:fusion_block compares the effect of different fusion schemes, i.e., static fusion versus dynamic fusion. In static fusion, two conventional linear layers are exploited to process the concatenated multi-modal features at each PoI. The dynamic fusion scheme refers to our proposed on in Section 3.4. Thanks to the adjustment of fusion layers’ parameters for each individual object query, our dynamic fusion outperforms the static setting by 1.4% NDS and 1.9% mAP.
Effect of Different Modalities.
In this experiment, we train models with single-modality input following the same setting as our default one. As shown in Table LABEL:tab:ablation:modality, the image-only baseline achieves 40.6% NDS and 29.4% mAP. Meanwhile, the LiDAR-only baseline achieves 70.2% NDS and 65.3% mAP, which is similar to TransFusion-L [2] (70.1% NDS and 65.1% mAP). Integrating the multi-modal information remarkably boosts the performance to $73.2\%$ NDS and $71.2\%$ mAP, validating the capability of our proposed fusion paradigm to integrate the complementary information from images and point clouds.
Effect of the Input Resolution.
The input resolution is also one of the key factors that influence the detection result. In Table LABEL:tab:ablation:img_resolution, we enlarge the image resolution from $800\times 448$ to $1600\times 640$, obtaining 0.5% NDS and 0.7% mAP gain. In Table LABEL:tab:ablation:voxel_size, we further analyze the influence of the voxel size. Either a larger image resolution or a smaller voxel size facilitates the feature representation of the corresponding modality, thus boosting the detection result. Our default setting employs an image resolution of $800\times 448$ and a voxel size of ($0.0075m,0.0075m,0.2m$), which is the same as most of recent works [2, 5, 67, 75].
Offsets (m)
0.0
0.2
0.4
0.6
0.8
1.0
NDS (%)
73.2
73.0
72.9
72.7
72.3
72.0
mAP (%)
71.2
71.1
70.8
70.5
70.0
69.4
Table 5: Sensor misalignment. Performance comparison on the impact of translation offsets due to calibration errors.
4.4 Detection Robustness Analysis
Sensor Misalignment.
In real applications, sensors can be misaligned due to physical impacts, installation errors, time-related drift, or some other unexpected reasons. This experiment investigates the robustness of PoIFusion against sensor misalignment. Specifically, we follow [2] to randomly add translation offsets to the calibration matrix at inference. The errors are formulated as uniformly distributed noise, with the maximum offset value varying from $0.2m$ to $1.0m$. As shown in Table 5, even when the maximum translation offset is $1.0m$, our method can achieve 72.0% NDS and 69.4% mAP, demonstrating the robustness of our fusion paradigm against sensor misalignment.
Sensor Failure.
In addition to sensor misalignment, we further explore a more serious condition, i.e., sensor failure. To validate our method under camera failure cases, we randomly drop several images by filling the corresponding images as all zeros. The results with different numbers of dropped images are presented in Table 6. The performance keeps dropping along with the increasing number of cameras that fail to work. Even with half of the cameras blocked, our PoIFusion achieves 68.6% mAP, obviously better than the LiDAR-only method. Moreover, in the extreme case that all of the cameras are not working, our method works comparable to the LiDAR-only baseline.
# Dropped Images
0
1
3
6
NDS (%)
73.2
72.8
72.0
70.2
mAP (%)
71.2
70.3
68.6
65.6
Discarded Sectors
${}^{\circ}$
${}^{\circ}$
${}^{\circ}$
${}^{\circ}$
${}^{\circ}$
NDS (%)
73.2
72.7
72.0
71.3
70.7
mAP (%)
71.2
70.2
68.9
67.6
66.4
Furthermore, to simulate the failure of a LiDAR sensor, we randomly discard point cloud data within a specified sectorial region based on the azimuthal angle. As shown in Table 7, in general, the impact of LiDAR failure is more pronounced than that of camera failure, since the spatial information captured by the LiDAR sensor is essential for precise localization. Notwithstanding the total absence of point cloud data within a $24^{\circ}$ sector, the mAP achieved by our method is still 1.1% higher than the LiDAR-only baseline. These two experiments demonstrate the robustness of our proposed PoIFusion against Sensor Failure.
5 Conclusion
In this work, we present PoIFusion, a query-based multi-modal 3D object detector, which is characterized by performing multi-modal feature sampling and fusion at generated points of interest. On the one hand, PoIFusion keeps the original view of the image and point cloud feature to preserve the modal-specific information. On the other hand, the employment of points of interest facilitates fine-grained fusion and enables representing object features in a flexible way. Our PoIFusion sets a state-of-the-art record on the highly competitive nuScenes dataset, and in the meantime keeps efficient runtime. Moreover, the experimental analysis demonstrates our method is robust to sensor misalignment and failure. We suppose our PoIFusion would serve as a strong baseline with huge potential for future exploration.
Supplementary Material
This supplementary material is composed of three sections. In Section A, we introduce supplementary details of our implementation. Then, in Section B, we provide additional experiments to further validate the merits of our proposed method. Finally, in section C, we visualize more examples to help intuitively present the detection results.
Appendix A Supplementary Details
Query Initialization.
As introduced in our main paper, we formulate each object query as a learnable dynamic 3D box $\bm{b}\in\mathbb{R}^{8}$ and an attached feature vector $\bm{v}\in\mathbb{R}^{256}$. Specifically, the query box is represented as:
$\bm{b}=[x_{c},y_{c},z_{c},w,l,h,sin\theta,cos\theta],$ (6)
where [$x_{c}$, $y_{c}$, $z_{c}$] is the center location, [$w$, $l$, $h$] is the box dimension, and $\theta$ is the azimuth angle. Notably, while velocity is a component of the detection results, it is not factored into the query boxes. When we initialize the query box, the BEV center location [$x_{c}$, $y_{c}$] is set to be uniformly distributed on the BEV, and the height center $z_{c}$ is set as 0. Besides, the dimensions $w$, $l$, $h$ of each box are respectively initialized as 6, 3, 2, and the azimuth is initialized as 0.
Distance-Biased Self-Attention Layer.
As outlined in Section 3.1, the multi-modal decoder utilizes a distance-biased self-attention layer [38] to capture the relationships and dependencies between different object queries. Here, we provide a comprehensive review of it.
Specifically, a conventional self-attention layer [56] calculates the attention score based on the dot product of query and key, which is detailed as:
[…]
where Q and K correspond to the query and key, and C is the channel dimension. The distance-biased self-attention layer incorporates a bias term that is contingent upon the centroid distance between pairs of query boxes. The adjusted attention score is computed as:
[…]
where $D$ indicates the centroid distance between query boxes, and $\psi$ is a scalar produced by a linear layer according to the query feature. In our multi-modal object decoder, we use 8 heads in the self-attention layer. The $\psi$ is not shared in different heads. We emphasize that the distance-biased self-attention layer is not the contribution of our work. We implement the self-attention layer with it since it improves the adaptability of the model with negligible costs.
Appendix B Additional Experimental Analysis
Different Weathers and Lighting Conditions.
The 3D object detection system on autonomous driving vehicles is supposed to be capable of working under varying lighting and weather conditions. In the main manuscript, we have shown two qualitative examples to give a glimpse of the robustness of our PoIFusion. In this experiment, we further follow [42] to divide the validation set of the nuScenes dataset into two pairs of subsets: sunny/rainy and day/night for comprehensive quantitative evaluation. The evaluated performance in terms of mAP (%) is summarized in Table S1. In general, the multi-modal 3D object detectors are not likely to be affected by the rainy weather. The changing of lighting condition significantly affects the detection performance, i.e., the mAP on the night subset is much lower than that of the day subset. Among the compared methods, our PoIFusion achieves the best performance under different weather and lighting conditions, further demonstrating the robustness of our fusion strategy.
@@ Appendix B Additional Experimental Analysis
Table S1: Different lighting and weather conditions. The reported performance is compared in terms of mAP (%) on the nuScenes validation set. We report the result of our PoIFusion with the Swin-T image backbone network.
Effect of Iteration Times of the Multi-Modal Decoder.
As introduced in Section 3.1, our multi-modal object decoder is iteratively applied to perform feature fusion and box refinement. In Table S2, we analyze how the iteration times of our multi-modal decoder influence the 3D object detection performance. Specifically, we vary the iteration times from 1 to 8 times and re-train the model for evaluation. In the beginning, the performance consistently improves along with the increasing iteration time. The peak performance is achieved when we apply the multi-modal object decoder 6 times. When we further increase the iteration time to 8, the performance drops for 0.5% NDS and 0.9% mAP. Therefore, in our default configuration, the multi-modal object decoder is iteratively applied 6 times. Moreover, as shown in the table, our method provides the potential to make a good trade-off between accuracy and latency. Specifically, when the iteration time is reduced to 4, our PoIFusion can still achieve the highest NDS (73.1%) and mAP (70.7%) among the compared methods presented in Table 1, and in the meantime run faster than the fastest competitor (i.e., BEVFusion) presented in Table 3.
# Iteration Times
NDS (%)
mAP (%)
Latency (ms)
1
67.1
63.4
87.7
2
71.9
69.3
94.2
4
73.1
70.7
99.8
6
73.2 71.2
115.2
8
72.7
70.3
119.2
Table S2: Effect of iteration times of the multi-modal decoder. The reported performance is compared on the nuScenes validation set. We exploit ResNet-50 as the image backbone network. The latency is evaluated on an NVIDIA A100 GPU.
Shared Parameters VS. Unshared Parameters
of the Multi-Modal Decoder. Since the multi-modal object decoder is iteratively applied multiple times, we can exploit either shared parameters or unshared parameters. The results of using shared parameters and unshared parameters are presented in Table S3. As shown in this comparison, these two models achieve comparable performance. Moreover, sharing parameters of the multi-modal object decoder makes a smaller model size. Therefore, we use shared parameters of the multi-modal decoder by default.
NDS (%)
mAP (%)
Shared Parameters
73.2
71.2
Unshared Parameters
73.3
70.9
Table S3: Shared versus unshared parameters of the multi-modal object decoder. The models are evaluated on the nuScenes validation set, with ResNet-50 as the image backbone network.
Training without the pretrained LiDAR encoder.
In our main experiment, we follow the training policy of most recent works [2, 42, 67, 5, 75], i.e., starting with the model weight of pretrained TransFusion-L and training the fusion model for 6 epochs. Here, we further study whether our PoIFusion can also be trained without a pretrained TransFusion-L model. Specifically, in this experiment, we follow the training policy of CMT [69] that trains the fusion framework with 20 epochs, with only the pretrained image backbone network. The performance comparison between our PoIFusion and CMT is presented in Table S4. Please note that these two models are trained with the same training policy, backbone network, and input resolution (details are summarized in the caption of this table). As shown in the table, our PoIFusion achieves 72.3% NDS and 68.4% mAP without pretrained LiDAR encoder, outperforming CMT with the same input resolution and training policy by 1.5% NDS and 0.5% mAP. This experiment demonstrates our PoIFusion works very well under various training policies, further exhibiting its potential to serve as a strong baseline for future research.
Method
NDS (%)
mAP (%)
CMT
70.8
67.9
PoIFusion (ours)
72.3 68.4
Table S4: Training without a pretrained TransFusion-L model. The compared models are trained with the same training policy and evaluated on the nuScenes validation set. Both models use ResNet-50 and VoxelNet as the backbone network, with image resolution set as $800\times 320$ and voxel size set as $(0.1m,0.1m,0.2m)$.
Appendix C Additional Qualitative Results
In our main manuscript, we only visualize two scenes due to the limitation of pages. Here, we add more qualitative examples in Figure S1 to help intuitively present the detection results of our method. From the visualized results, it can be seen that our proposed method is able to accurately detect 3D objects under complex road conditions.
Figure S1: Supplement of Qualitative results on the nuScenes validation set.

### v2

@@ 4.3.2 Results on Argoverse2
Table 3: Performance comparison of 3D multi-object tracking on nuScenes validation split in terms of AMOTA (%) and AMOTP (%).
Methods AMOTA $\uparrow$ AMOTP $\downarrow$
CenterPoint Yin et al. (2021a) 63.7 60.6
VoxelNeXt Chen et al. (2023a) 70.2 64.0
TransFusion Bai et al. (2022) 71.8 60.3
BEVFusion (MIT) Liu et al. (2023b) 72.8 59.4
ObjectFusion Cai et al. (2023) 74.2 54.3
PoIFusion(ours) 75.1 50.7
4.3.3 Extended Results for Multiple Object Tracking
In addition to the main results of 3D object detection, we evaluate the generalizability of our method on the 3D multiple object tracking task of the nuScenes dataset. Specifically, we follow CenterPoint Yin et al. (2021a) to adopt the “tracking-by-detection” scheme that offline links the detection boxes into tracking tubelets and to evaluate the performance in terms of AMOTA and AMOTP. We would like to clarify that the compared methods, except VoxelNeXt Chen et al. (2023a), all follow the same “tracking-by-detection” scheme. As shown in Table 3, our PoIFusion works the best (75.1% AMOTA and 50.7% AMOTP) among the compared methods. This experiment shows that our better detection results can facilitate the downstream task, i.e., 3D multiple object tracking, which is also an important component in the perception system of an autonomous-driving vehicle.
@@ 4.3.3 Extended Results for Multiple Object Tracking
Table 4: Latency comparison. The latency is evaluated on nuScene, with batch size set as 1. $\dagger$”: accelerated with Flash-Attention Dao et al. (2022) (NVIDIA V100 GPU is not compatible with the Flash Attention operator). Top-2 entries are with bold font.
Methods Latency (ms)
A100 V100
TransFusion Bai et al. (2022) 153.8 190.3
BEVFusion (MIT) Liu et al. (2023b) 113.9 135.5
DeepInteraction Yang et al. (2022c) 204.1 344.8
CMT Yan et al. (2023) 210.8 387.1
CMT Yan et al. (2023) 159.7 -
PoIFusion (ours) 115.2 163.7
4.3.4 Latency Comparison
Furthermore, we conduct the latency comparison among the recent works. For a fair comparison, we re-evaluate the latency on an NVIDIA A100 and an NVIDIA V100 GPU. Since the implementation of the voxelization operation will not influence the detection results but becomes a confounder for runtime evaluation, we exclude the costed time of voxelization in this comparison. As shown in Table 4, the latency of our PoIFusion is 115.2ms on an NVIDIA A100 GPU, which is comparable to the fastest method, i.e., BEVFusion (MIT), which takes 113.9ms to process each sample. This comparison further demonstrates the potential of our proposed PoIFusion to serve as a strong baseline for future investigation and application.
Table 5: Ablative experiments of components in the PoI multi-modal decoder. The backbone networks are ResNet-50 and VoxelNet. Image resolution is set as $800\times 448$, and the voxel size of point clouds is ($0.075m$,$0.075m$,$0.2m$). Our default setting is highlighted in gray.
Anchor points NDS (%) mAP (%)
Center only 72.8 70.6
Center + corner 73.2 71.2
(a) The choice of anchor points. Exploiting both the center and corner points as anchor points for points of interest generation yields better performance.
PoI generation NDS (%) mAP (%)
Baseline 72.4 70.1
+ B.T. 72.8 70.8
+ B.T. & P.S. 73.2 71.2
(b) Operations to derive PoIs. The baseline in this experiment directly uses anchor points as PoIs. B.T.: box transformation. P.S.: point shift.
Fusion block NDS (%) mAP (%)
Static fusion 71.8 69.3
Dynamic fusion 73.2 71.2
(c) Fusion block. Dynamic fusion block produces adaptive parameters for the fusion layer, improving the performance.
4.4 Experimental Analysis
In this section, we conduct several groups of controlled experiments to experimentally analyze the effects of each factor in our proposed PoIFusion framework. In these experiments, the image backbone network is ResNet-50 and the point cloud backbone network is VoxelNet. Unless specified, the image resolution is set as $800\times 448$ and the voxel size is set as ($0.075m$,$0.075m$,$0.2m$). The models are all evaluated on the validation set of the nuScenes dataset.
4.4.1 Components of PoI Multi-Modal Decoder
In this section, we provide a detailed analysis of the key components in the PoI multi-modal decoder. The experiments are structured to investigate three critical aspects: (a) the selection of anchor points, (b) the generation of PoIs from anchor points, and (c) the fusion strategy to integrate multi-modal information.
As shown in Table 5, the model that only uses the center point as the anchor to generate PoIs achieves 72.8% NDS and 70.6% mAP. By additionally involving the corner points as the anchor points, the NDS and mAP boosts to 73.2% and 71.2%, respectively. This comparison shows the benefits of representing the geometric property of query boxes.
In Table 5, we validate the operations to derive PoIs from anchor points. The baseline directly takes the anchor points as PoIs without adaptation, achieving 70.1% mAP. By involving the holistic box-level transformation and further applying the point-wise shift, the performance boosts to 70.8 % mAP and 71.2% mAP, respectively. The online adjustment of PoIs ameliorates the misalignment in feature sampling and thereby boosts the detection result.
Table 5 compares the effect of different fusion schemes, i.e., static fusion versus dynamic fusion. In static fusion, two conventional linear layers are exploited to process the concatenated multi-modal features at each PoI. The dynamic fusion scheme refers to our proposed on in Section 3.5. Thanks to the adjustment of fusion layers’ parameters for each object query, our dynamic fusion outperforms the static setting by 1.4% NDS and 1.9% mAP.
Table 6: Effect of leveraging different modalities. Integrating RGB images and LiDAR point clouds remarkably boosts the detection result.
Modality NDS (%) mAP (%)
Camera only 40.6 29.4
LiDAR only 70.2 65.3
Camera + LiDAR 73.2 71.2
4.4.2 Effect of Different Modalities
In this experiment, we train models with single-modality input following the same training setting as introduced in Section 4.2. This experiment is presented in Table 6. The image-only baseline achieves 40.6% NDS and 29.4% mAP, and the LiDAR-only baseline achieves 70.2% NDS and 65.3% mAP. Integrating the multi-modal information remarkably boosts the performance to $73.2\%$ NDS and $71.2\%$ mAP. Please note that the performance of our LiDAR-only baseline is comparable to that of TransFusion-L (70.1% NDS and 65.1% mAP), which is the LiDAR-only baseline of Bai et al. (2022); Liang et al. (2022); Liu et al. (2023b); Yang et al. (2022c); Xie et al. (2023); Cai et al. (2023), but our fusion model makes a more significant improvement. This comparison further validates the benefits of our proposed fusion at points of interest scheme to integrate complementary information from images and point clouds.
4.4.3 Effect of the Input Resolution
The input resolution (i.e., image resolution, and voxel size) is also one of the key factors that influence the detection result. In this experiment, we present the performance comparison under various input resolutions.
Image Resolution. As shown in Table 7, we enlarge the image resolution from $800\times 448$ to $1600\times 640$, which is the same as the input resolution of the strongest competitor CMT. Compared to our default setting, the detection performance is observed with an improvement of 0.5% NDS and 0.7% mAP. The larger resolution provides more detailed textual information from the RGB images and thus benefits multi-modal 3D object detection. It is worth noting that although increase the image resolution can definitely improve the performance, we have not increase it in our default setting, which ensures fair comparison with others.
Table 7: Ablative experiments of the input resolution. The backbone networks are ResNet-50 and VoxelNet. Our default setting is highlighted in gray.
Image resolution NDS (%) mAP (%)
$800\times 448$ 73.2 71.2
$1600\times 640$ 73.7 71.9
(a) Image Resolution. Increasing the image resolution strengthens the textual information provided by images.
Voxel size NDS (%) mAP (%)
(0.125, 0.125, 0.2) 71.6 69.1
(0.1, 0.1, 0.2) 72.3 69.8
(0.075, 0.075, 0.2) 73.2 71.2
(b) Voxel Size. A smaller voxel size corresponds to a larger resolution of the point cloud feature, facilitating the localization information for 3D object detection.
Voxel Size. In Table 7, we further analyze the influence of the voxel size by setting the voxel size as (0.125m, 0.125m, 0.2m), (0.1m, 0.1m, 0.2m) and (0.075m, 0.075m, 0.2m). As shown in the table, the model with a smaller voxel size works better than that with a larger one. Smaller voxel size results in a larger resolution of the point cloud feature map, which facilitates more precise localization information.
@@ 4.4.3 Effect of the Input Resolution
Table 8: Performance comparison between using shared parameters and unshared parameters for the PoI multi-modal decoder. Our default setting is highlighted in gray.
Parameters NDS (%) mAP (%)
Unshared parameters 73.3 70.9
Shared parameters 73.2 71.2
4.4.4 Shared Parameters VS. Unshared Parameters for the PoI Multi-Modal Decoder
Our PoI multi-modal decoder is applied iteratively, with the option to use either shared or unshared parameters. The performance comparison between these two configurations is shown in Table 8, where both models demonstrate comparable results. However, employing shared parameters reduces the overall model size, making it more efficient. Consequently, we adopt the shared-parameter configuration as the default setting for the PoI multi-modal decoder.
Table 9: Different weathers and lighting conditions. The reported performance is compared in terms of mAP (%) on the nuScenes validation set. We report the result of our PoIFusion with the Swin-T image backbone network, which is consistent with MVP Yin et al. (2021b) and BEVFusion Liu et al. (2023b).
@@
4.5 Robustness Analysis
Besides the accuracy and efficiency, the capability of working under different environments and corruptions is also what we expected for a desirable perception module in autonomous driving vehicles. In this section, we present a series of experiments to examine the robustness of our PoIFusion framework.
4.5.1 Different Weathers and Lighting Conditions
The 3D object detection system on autonomous driving vehicles is supposed to be capable of working under varying lighting and weather conditions. In this experiment, we follow Liu et al. (2023b) to divide the validation set of the nuScenes dataset into two pairs of subsets: sunny/rainy and day/night for comprehensive quantitative evaluation. The evaluated performance in terms of mAP (%) is summarized in Table 9. In general, the multi-modal 3D object detectors are not likely to be affected by the rainy weather. The changing of lighting condition significantly affects the detection performance, i.e., the mAP on the night subset is much lower than that of the day subset. Among the compared methods, our PoIFusion achieves the best performance under different weather and lighting conditions, validating the robustness of fusion at points of interest.
Moreover, we visualize the qualitative results of detecting 3D objects under different weather and lighting conditions in Figure 4. We present one example from a sunny day and another example from a rainy night. These examples further give an intuitive understanding that PoIFusion can robustly perform 3D object detection under varying weather and lighting conditions
@@ 4.5.1 Different Weathers and Lighting Conditions
Table 10: Sensor misalignment. Performance comparison on the impact of translation offsets due to calibration errors. The reported performance is compared in terms of mAP (%) on the nuScenes validation set. We report the result of our PoIFusion with the ResNet-50 image backbone network.
Misalignment offset (m) NDS (%) mAP (%)
0.0 73.2 71.2
0.2 73.0 71.1
0.4 72.9 70.8
0.6 72.7 70.5
0.8 72.3 70.0
1.0 72.0 69.4
4.5.2 Robustness Against Sensor Misalignment
In real applications, sensors can be misaligned due to physical impacts, installation errors, time-related drift, or some other unexpected reasons. This experiment investigates the robustness of PoIFusion against sensor misalignment. Specifically, we follow Bai et al. (2022) to randomly add translation offsets to the calibration matrix at inference. The errors are formulated as uniformly distributed noise, with the maximum offset value varying from $0.2m$ to $1.0m$. As shown in Table 10, even when the maximum translation offset is $1.0m$, our method can achieve 72.0% NDS and 69.4% mAP, still improving our LiDAR-only baseline with 1.8% NDS and 4.1% mAP. This experiment demonstrates the robustness of our fusion paradigm against sensor misalignment.
4.5.3 Robustness Against Sensor Failure
In addition to sensor misalignment, we further explore a more terrible corruption, i.e., sensor failure.
Camera Failure. To validate our method under camera failure cases, we randomly drop several images by filling the corresponding images as all zeros. The results with different numbers of dropped images are presented in Table 11. The performance keeps dropping along with the increasing number of cameras that fail to work. Even with half of the cameras blocked, our PoIFusion achieves 68.6% mAP, obviously better than the LiDAR-only method. Moreover, in the extreme case that all of the cameras are not working, our method still works a little bit better than the LiDAR-only baseline, showing the benefit of multi-modal joint training.
# Dropped Images NDS (%) mAP (%)
0 73.2 71.2
1 72.8 70.3
3 72.0 68.6
6 70.2 65.6
Discarded Sectors NDS (%) mAP (%)
$0^{\circ}$ 73.2 71.2
$6^{\circ}$ 72.7 70.2
$12^{\circ}$ 72.0 68.9
$18^{\circ}$ 71.3 67.6
$24^{\circ}$ 70.7 66.4
LiDAR Failure. Furthermore, to simulate the failure of a LiDAR sensor, we randomly discard point cloud data within a specified sectorial region based on the azimuthal angle. As shown in Table 12, in general, the impact of LiDAR failure is more pronounced than that of camera failure, since the information captured by the LiDAR sensor is essential for precise localization. Notwithstanding the total absence of point cloud data within a $24^{\circ}$ sector, the mAP achieved by our method is still 1.1% higher than the LiDAR-only baseline. These two experiments demonstrate the robustness of our proposed PoIFusion against Sensor Failure.
5 Conclusion
In this work, we introduce PoIFusion, a query-based multi-modal 3D object detector that leverages multi-modal feature sampling and fusion at strategically generated points of interest. PoIFusion preserves modal-specific information by maintaining the original views of both image and point cloud features, while the use of points of interest allows for fine-grained fusion and a flexible representation of object features. Our approach establishes a new state-of-the-art on the challenging nuScenes dataset, all while maintaining efficient runtime performance. Furthermore, experimental results highlight the robustness of PoIFusion to sensor misalignment and failure. We believe that PoIFusion will serve as a robust baseline with significant potential for future research and deployment in the field.
6 Data Availability
The nuScenes dataset Caesar et al. (2020) and Argoverse2 dataset Wilson et al. (2021) used in this study are well-recognized benchmarks which are public available. Specifically, nuScenes dataset can be accessed through https://www.nuscenes.org/ and Argoverse2 dataset can be accessed through https://www.argoverse.org/av2.html. We have not used extra private data in our experiments.